OpenStudy (anonymous):

Find the eigenvalues and corresponding set of basic eigenvectors for -3 2 -4 3

OpenStudy (anonymous):

you mean you have eigenvector now backward to get eigenvalue?

OpenStudy (anonymous):

and yours is\[\left[\begin{matrix}-3 & 2 \\ -4 & 3\end{matrix}\right] or \]\[\left[\begin{matrix}-3 & -4 \\ 2 & 3\end{matrix}\right] \]?

OpenStudy (anonymous):

ridiculous asker!!!